OpenStudy (anonymous):

HELPS ASAPP!!!! MEDAL +FAN!! On which of the following lines does the point (7,1) lie? A. y - 5x = 4 B. y= x - 10 C. y= -4x + 3 D. y=-x + 8

OpenStudy (kidrah69):

try plugging in the coordinate to each equation

OpenStudy (anonymous):

oh ok give me a couple min

OpenStudy (kidrah69):

1 \[1-5(7)=4\] 2\[1=7-10\] 3\[1=-4(7)+3\] 4\[1=-7+8\]

OpenStudy (anonymous):

Thanks! i will complete them

OpenStudy (kidrah69):

yea and then just see which statement is true and you have your answer! :)

OpenStudy (anonymous):

i thought it said lie like the false one

OpenStudy (kidrah69):

that cant be right...theres only one of the statements that is true there are 3 false statements

OpenStudy (anonymous):

i will see and then tag you or if you want to wait for what i get :)

OpenStudy (kidrah69):

ill wait :) let me know if u need help :3

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

D is true and the rest are false so lie = true?

OpenStudy (kidrah69):

i think lie means where does the point reside i agree with D :)

OpenStudy (kidrah69):

good job :)

OpenStudy (anonymous):

Thanks again! :D

OpenStudy (kidrah69):

no probs :)